A pair of fair dice is rolled. Let X the random variable representing the sum of the numbers that appear.Find P(3 ≤X≤ 10)

1
Expert's answer
2022-02-22T13:49:37-0500

The sample space is,

"S=\begin{Bmatrix}
 (1,1) & (2,1)&(3,1)&(4,1)&(5,1)&(6,1) \\
 (1,2) & (2,2)&(3,2)&(4,2)&(5,2)&(6,2)\\
(1,3)&(2,3)&(3,3)&(4,3)&(5,3)&(6,3)\\
(1,4)&(2,4)&(3,4)&(4,4)&(5,4)&(6,4)\\
(1,5)&(2,5)&(3,5)&(4,5)&(5,5)&(6,5)\\
(1,6)&(2,6)&(3,6)&(4,6)&(5,6)&(6,6)
\end{Bmatrix}"

This sample space shows the outcome on the first dice followed by the outcome on the second dice.


Let the random variable "Z" represent the sum of the outcome on the first dice and the outcome on the second dice.

Taking the sum of the outcome on the first and second dice gives,

"Z=\begin{Bmatrix}
 2 & 3&4&5&6&7 \\
 3 & 4&5&6&7&8\\
4&5&6&7&8&9\\
5&6&7&8&9&10\\
6&7&8&9&10&11\\
7&8&9&10&11&12

\end{Bmatrix}"

Clearly, the random variable "Z" may take on the values 2,3,4,5,6,7,8,9,10,11,12 with its probability distribution given as,


"z" 2 3 4 5 6 7 8 9 10 11 12

"p(z)" "1\over36" "2\over36" "3\over36" "4\over36" "5\over36" "6\over36" "5\over36" "4\over36" "3\over36" "2\over36" "1\over36"

Now,


"P(3 \le X\le10)={2\over36}+{3\over36}+{4\over36}+{5\over36}+{6\over36}+{5\over36}+{4\over36}+{3\over36}={32\over36}={8\over9}"

The required probability is "{8\over9}".